# Two-dimensional FrBD friction models for rolling contact: extension to linear viscoelasticity

### v4

@@ 5 Numerical simulations
In the following, FrBD2-GM and FrBD3-GM Models 3.2, 3.2, and Models 3.2 are numerically simulated and compared to the corresponding FrBD1-KV versions concerning both their steady-state and transient behaviours. More specifically, Sect. 5.1 is dedicated to steady rolling contact, whereas unstationary phenomena are investigated in Sect. 5.2. As discussed in Appendix B, the FrBD2-GM models correspond directly to a Standard Linear Solid element, for which the interconversion to an equivalent FrBD2-KV parametrisations is available analytically. In the remainder of the paper, the GM formulation is preferred because it is more widely used in the literature, and also because yields a more intuitive state-space reformulation for the FrBD2-SLS dynamics (see Appendix B). Additional implementational and computational details may be found in [88].
@@ 5.1 Steady-state behaviour
Due to the distributed nature of the rolling contact process, the relaxation phenomena introduced by viscoelastic components in higher-order rheological representations of the bristle element are not limited to transient regimes, but manifest even under steady-state conditions. Indeed, as suggested already by Eq. (54), structural damping acts also spatially, contributing to energy dissipation as the system’s dynamics evolves along the rolling direction. It is therefore of interest to investigate how more refined viscoelastic descriptions of the bristle dynamics affect the so-called force-slip surfaces, namely the mapping $(\hat{\bm{F}}_{\bm{x}},\hat{M}_{z})\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{4}\mapsto\mathbb{R}^{3}$ given by
@@
In the absence of spin ($\bm{\varphi}=\bm{0}$), the main characteristics obtained for rectangular and elliptical contact areas with parabolic pressure distributions, and using the model parameters listed in Table 1, are shown in Figs. 4 and 5. Assuming isotropic friction, the matrix $\mathbf{M}(\bm{v}_{\textnormal{s}})$ was modelled as $\mathbf{M}(\bm{v}_{\textnormal{s}})=\mu(\bm{v}_{\textnormal{s}})\mathbf{I}_{2}$, with
@@
where $\mu_{\textnormal{s}},\mu_{\textnormal{d}}\in\mathbb{R}_{>0}$ denote the static and dynamic friction coefficients, $v_{\textnormal{S}}\in\mathbb{R}_{>0}$ the Stribeck velocity, $\delta_{\textnormal{S}}\in\mathbb{R}_{\geq 0}$ the Stribeck exponent, and $\mu_{\textnormal{v}}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{2}\mapsto\mathbb{R}_{\geq 0}$ captures the viscous friction.
Compared with the FrBD1-KV formulation employed in [88], inspection of Figs. 4 and 5 indicates that the additional relaxation effects introduced by the FrBD2-GM and FrBD3-GM models tend to attenuate both the tangential forces and moments, yielding trends of slightly lower magnitude. The effect is particularly evident in the vertical moment, for both rectangular and elliptical contact areas. These conclusions remain valid when considering the presence of large spin slips, as illustrated in Figs. 4 and 5.
@@ 5.2 Transient behaviour
First, the system’s response to a step input in the rigid relative velocity is considered. Figure 8, produced using Models 3.2 in line contact, illustrates the tangential forces and vertical moment for a step slip input $\bm{\sigma}=(0.08,0.16)$. The FrBD1-KV model (solid thick lines) exhibits a response closely resembling a first-order system in all characteristics. In contrast, the higher-order FrBD2-GM and FrBD3-GM models (solid and dashed lines) display richer dynamics, including an initial overshoot characteristic of nonlinear systems. Notably, whilst the FrBD1-KV formulation predicts convergence of the forces and moment to their steady-state values for $s\leq 0.1$ m, the refined models require approximately $s=0.3$ m. This highlights the significant role of hysteretic damping in shaping the transient response of rolling contact elements, including pneumatic and non-pneumatic tyres.
@@
Figures 9 and 10 visualise the tangential forces and vertical moment for $\bar{\bm{\sigma}}=(0.04,0.08)$ and $(\omega_{1},\omega_{2})=(50,0)$, $(\omega_{1},\omega_{2})=(50,100)$ $\textnormal{m}^{-1}$, respectively. The trends confirm the previous analysis: whilst the simplest model reproduces the input oscillations almost immediately, the more sophisticated FrBD2-GM and FrBD3-GM introduce additional delay effects that are not adequately captured by standard rheological representations. This discrepancy is particularly pronounced for the vertical moment, which is consistent with the observations reported in Sect. 5.1.
@@ 6 Conclusions
Finally, it should be emphasised that the models presented in this paper are exclusively based on classical derivative rheological descriptions. Recent studies have shown that fractional-derivative models can successfully capture viscoelasticity with fewer parameters. Future work may therefore focus on deriving more sophisticated formulations that combine PDEs with fractional ODEs, providing physically and thermodynamically consistent descriptions of viscoelastic rolling contact.
@@ A.1 Linear viscoelastic response in the frequency domain
where $\mathrm{Sym}_{2}(\mathbb{R})\ni\mathbf{K}_{i}\succ\mathbf{0}$, $i\in\{0,\dots,n\}$, and $\mathrm{Sym}_{2}(\mathbb{R})\ni\mathbf{C}_{i}\succ\mathbf{0}$, $i\in\{1,\dots,n\}$, denote stiffness and damping matrices, i.e., $\mathbf{K}_{i}=p\bar{\mathbf{K}}_{i}$, and $\mathbf{C}_{i}=p\bar{\mathbf{C}}_{i}$. An equivalent expression can be derived for the GKV representation, yielding the same complex modulus $\mathbf{G}^{*}(\omega)$, in accordance with the equivalence stated in Eq. (10). Equation (69) highlights that GM and GKV elements naturally generate a discrete relaxation spectrum, with characteristic relaxation time matrices $\bm{\tau}_{i}=\bar{\mathbf{C}}_{i}\bar{\mathbf{K}}_{i}^{-1}=\mathbf{C}_{i}\mathbf{K}_{i}^{-1}$, $i\in\{1,\dots,n\}$.
@@ A.3 FrBDn+1 model parameters and interconversions
Once the stiffness and damping matrices $\mathbf{K}_{i}$ and $\mathbf{C}i$ have been identified, they can be directly employed in the FrBDn+1{n+1} models introduced in Sect. 2.2 by introducing the normalised parameters $\bar{\mathbf{K}}_{i}=\mathbf{K}_{i}/p$, $i\in\{0,\dots,n\}$, and $\bar{\mathbf{C}}_{i}=\mathbf{C}_{i}/p$, $i\in\{1,\dots,n\}$. where $p(\bm{x},s)$ denotes the local normal pressure. For the GM realisation, these parameters enter the dynamics through Eq. (19), whereas for the GKV realisation they appear in Eq. (20).

### v5

# arXiv:2601.13818v5[physics.app-ph] 02 Mar 2026

@@ 5 Numerical simulations
In the following, FrBD2-GM and FrBD3-GM Models 3.2, 3.2, and Models 3.2 are numerically simulated and compared to the corresponding FrBD1-KV versions concerning both their steady-state and transient behaviours. More specifically, Sect. 5.1 is dedicated to steady rolling contact, whereas unstationary phenomena are investigated in Sect. 5.2. As discussed in Appendix B, the FrBD2-GM models correspond directly to a Standard Linear Solid element, for which the interconversion to an equivalent FrBD2-KV parametrisation is available analytically. In the remainder of the paper, the GM formulation is preferred because it is more widely used in the literature, and also because yields a more intuitive state-space reformulation for the FrBD2-SLS dynamics (see Appendix B). Additional implementational and computational details may be found in [88].
@@ 5.1 Steady-state behaviour
For lumped parameter friction models, all the FrBDn+1 variants collapse to the same force-slip relationship given by Eq. (21). This confirms analytically that increasing the rheological order does not alter the stationary creep characteristics: the asymptotic friction level is determined solely by the zeroth-order branch stiffness. Physically, this reflects the fact that, under constant sliding, all viscoelastic branches reach equilibrium and no internal stress relaxation remains active. Due to the distributed nature of the rolling contact process, however, the relaxation phenomena introduced by viscoelastic components in higher-order, distributed parameter rheological representations of the bristle element are not limited to transient regimes, but manifest even under steady-state conditions. Indeed, as suggested already by Eq. (54), structural damping acts also spatially, contributing to energy dissipation as the system’s dynamics evolves along the rolling direction. It is therefore of interest to investigate how more refined viscoelastic descriptions of the bristle dynamics affect the so-called force-slip surfaces, namely the mapping $(\hat{\bm{F}}_{\bm{x}},\hat{M}_{z})\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{4}\mapsto\mathbb{R}^{3}$ given by
@@
In the absence of spin ($\bm{\varphi}=\bm{0}$), the main characteristics obtained for rectangular and elliptical contact areas with parabolic pressure distributions, and using the model parameters listed in Table 1, are shown in Figs. 4 and 5. The geometrical and rheological parameters reported in Table 1 are of the same order of magnitude as those adopted in previous studies on rubber friction (e.g., [78, 79, 81, 82]) and are consistent with standard values commonly used in the tyre dynamics literature [8]. In particular, the longitudinal and lateral micro-stiffnesses associated with the higher-order branches were selected to be close to those of the zeroth branch. This choice ensures a comparable steady-state response across all model variants, thereby allowing for a clearer assessment of the differences in viscoelastic behaviour arising specifically from the damping and relaxation coefficients. Assuming isotropic friction, the matrix $\mathbf{M}(\bm{v}_{\textnormal{s}})$ was modelled as $\mathbf{M}(\bm{v}_{\textnormal{s}})=\mu(\bm{v}_{\textnormal{s}})\mathbf{I}_{2}$, with
@@
where $\mu_{\textnormal{s}},\mu_{\textnormal{d}}\in\mathbb{R}_{>0}$ denote the static and dynamic friction coefficients, $v_{\textnormal{S}}\in\mathbb{R}_{>0}$ the Stribeck velocity, $\delta_{\textnormal{S}}\in\mathbb{R}_{\geq 0}$ the Stribeck exponent, and $\mu_{\textnormal{v}}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{2}\mapsto\mathbb{R}_{\geq 0}$ captures the viscous friction. Also in this case, the values of the friction parameters listed in Table 1 were adapted from [78, 79, 81, 82].
The trends illustrated in Figs. 4 and 5 are in good qualitative agreement with the force surfaces predicted by other physical-analytical and empirical models of tyre dynamics, such as standard brush models with generalised Coulomb friction and Pacejka’s Magic Formula. In particular, adopting a sliding-velocity-dependent friction coefficient as in Eq. (66) permits reproducing the characteristic force peak at low slip, followed by a progressive decrease in tangential force toward an asymptotic value at large slips. In this respect, it is worth noting that, whilst LuGre-like formulations of rolling contact friction correctly postulate the friction coefficient as a function of the sliding velocity [78, 79, 81, 82], Pacejka’s Magic Formula and several modified brush models typically assume a slip-dependent equation. As a consequence, the resulting force-slip surfaces are inherently independent of the rolling velocity. However, empirical evidence suggests that the tangential forces and vertical moment do depend on rolling speed [78], as predicted by sliding-velocity-dependent friction laws similar to Eq. (66). Importantly, the capability to consistently combine higher-order viscoelastic rheological behaviour with sliding-velocity-dependent friction is essential in rubber friction applications [108, 109, 110]. In such cases, slow relaxation dynamics may induce local heating and trigger flash-temperature effects, which in turn modify the frictional response [109, 110]. Accounting for these coupled mechanisms is therefore crucial for a physically sound and predictive modelling framework.
Local hysteresis affects rolling contact friction also directly: compared with the FrBD1-KV formulation employed in [88], inspection of Figs. 4 and 5 indicates that the additional relaxation effects introduced by the FrBD2-GM and FrBD3-GM models tend to attenuate both the tangential forces and moments, yielding trends of slightly lower magnitude. Phenomenologically, this may again be explained by observing that, due to the distributed character of the rolling contact process, viscous damping acts spatially along the contact patch length, producing smaller bristle deformations and forces in the proximity of the trailing edge, as also confirmed in simulation. The attenuation is particularly evident in the vertical moment, for both rectangular and elliptical contact areas, where the presence of relaxation over multiple time scales appears to influence not only its magnitude, but also the position of its peak value. These conclusions remain valid when considering the presence of large spin slips, as illustrated in Figs. 6 and 7. In this context, particular attention should be given to the role of viscoelastic relaxation in mitigating the asymmetrisation effects induced by large effective spin values, as observed in both Figs. 6(b) and 7(b). This behaviour is physically intuitive and arises from the same mechanisms discussed above: viscous damping attenuates nonlinear coupling interactions by reducing the bristle deformations, which, in turn, results in a more symmetric trend of the lateral and longitudinal forces.
Additional numerical investigations, not reported here for the sake of brevity, were carried out by varying the rheological parameters listed in Table 1 whilst keeping $n=1$ and $n=2$ fixed. These parametric studies did not reveal any substantial differences in the predicted response, indicating that the model behaviour is essentially determined by the rheological order, but relatively insensitive to moderate variations of the adopted coefficients.
@@ 5.2 Transient behaviour
First, the system’s response to a step input in the rigid relative velocity is considered. Figure 8, produced using Models 3.2 in line contact, illustrates the tangential forces and vertical moment for a step slip input $\bm{\sigma}=(0.08,0.16)$. The FrBD1-KV model (solid thick lines) exhibits a response closely resembling a first-order system in all characteristics. In contrast, the higher-order FrBD2-GM and FrBD3-GM models (solid and dashed lines) display richer dynamics, including an initial overshoot characteristic of nonlinear systems. Notably, whilst the FrBD1-KV formulation predicts convergence of the forces and moment to their steady-state values for $s\leq 0.1$ m, the refined models require approximately $s=0.3$ m. This highlights the significant role of hysteretic damping in shaping the transient response of rolling contact elements, including pneumatic and non-pneumatic tyres. An interpretation of the behaviour depicted in Fig. 8 may be obtained by observing that, in the distributed formulation, time relaxation is convected spatially along the contact patch. For a constant rolling speed $V_{\textnormal{r}}(s)=V_{\textnormal{r}}$, each relaxation time $\bm{\tau}_{i}$, $i\in\{1,\dots,n\}$, corresponds to a characteristic relaxation length. When the imposed slip variation occurs over distances comparable to or smaller than these relaxation lengths, the dissipative branches cannot fully equilibrate before exiting the contact region. As a consequence, elastic energy is temporarily stored in the slower branches and released downstream, producing the observed overshoot and delayed convergence. The FrBD1-KV model, being associated with a single relaxation time, can reproduce only one characteristic length scale and therefore cannot capture the multi-stage adjustment process emerging in the FrBD2 and FrBD3 formulations.
@@
Figures 9 and 10 visualise the tangential forces and vertical moment for $\bar{\bm{\sigma}}=(0.04,0.08)$ and $(\omega_{1},\omega_{2})=(50,0)$, $(\omega_{1},\omega_{2})=(50,100)$ $\textnormal{m}^{-1}$, respectively. The trends confirm the previous analysis: whilst the simplest model reproduces the input oscillations almost immediately, the more sophisticated FrBD2-GM and FrBD3-GM introduce additional delay effects that are not adequately captured by standard rheological representations. This discrepancy is particularly pronounced for the vertical moment, which is consistent with the observations reported in Sect. 5.1. From a frequency-domain viewpoint, the observed differences may be explained by recalling that a GM element exhibits a storage and loss modulus that vary with excitation frequency (see Appendix A). The sinusoidal slip input, therefore, probes the effective complex modulus of the bristle element. Whilst the FrBD1-KV model behaves essentially as a single-pole low-pass filter, the higher-order GM representations introduce multiple poles associated with distinct relaxation times, resulting in amplitude attenuation and phase lag that depend on the spectral content of the imposed slip variation.
The phenomena discussed above have profound repercussions on the dynamics of road vehicles. Indeed, they imply that, in tyres, rapid manoeuvres or high-frequency excitation – such as those encountered during braking on rough surfaces or wheel shimmy phenomena – may activate viscoelastic mechanisms that remain essentially invisible in first-order models. More generally, the rheological order becomes particularly relevant when the characteristic time scales of vehicle manoeuvres are comparable to the intrinsic material relaxation times. Under slow quasi-static conditions, all formulations collapse to nearly identical responses; however, in highly dynamic scenarios, the simplified FrBD1-KV model may underestimate transient peaks in forces and moments. Since such peaks are known to influence stability margins in lateral dynamics and braking performance, incorporating experimentally identified relaxation spectra may improve predictive accuracy without altering the underlying friction law.
@@ 6 Conclusions
Compared to classical hereditary viscoelastic rolling formulations, which involve time convolution integrals, the approach presented in this paper provides a state-space realisation with explicit internal variables, facilitating coupling with multibody simulations and control algorithms while preserving consistency with linear viscoelastic material theory. In particular, in contrast to phenomenological tyre brush models, where relaxation effects are embedded in empirical relaxation lengths, the FrBDn+1 framework links transient rolling behaviour directly to experimentally identifiable material spectra. Finally, it should be emphasised that the models developed in this manuscript are exclusively based on classical derivative rheological descriptions. Recent studies have shown that fractional-derivative models can successfully capture viscoelasticity with fewer parameters. Future work may therefore focus on deriving more sophisticated formulations that combine PDEs with fractional ODEs, providing physically and thermodynamically consistent descriptions of viscoelastic rolling contact.
@@
  • [108] Persson BNJ. Theory of rubber friction and contact mechanics. The Journal of Chemical Physics 115(8):3840-3861 (2001). Available from: https://doi.org/10.1063/1.1388626.
  • [109] Persson BNJ. Rubber friction: role of the flash temperature. J. Phys.: Condens. Matter 18 7789 (2006). Available from: https://doi.org/10.1088/0953-8984/18/32/025.
  • [110] Mavros G. A thermo-frictional tyre model including the effect of flash temperature. Vehicle System Dynamics, 57(5):721-751 (2019). Available from: https://doi.org/10.1080/00423114.2018.1484147.
@@ A.1 Linear viscoelastic response in the frequency domain
where $\mathbf{Sym}_{2}(\mathbb{R})\ni\mathbf{K}_{i}\succ\mathbf{0}$, $i\in\{0,\dots,n\}$, and $\mathbf{Sym}_{2}(\mathbb{R})\ni\mathbf{C}_{i}\succ\mathbf{0}$, $i\in\{1,\dots,n\}$, denote stiffness and damping matrices, i.e., $\mathbf{K}_{i}=p\bar{\mathbf{K}}_{i}$, and $\mathbf{C}_{i}=p\bar{\mathbf{C}}_{i}$. An equivalent expression can be derived for the GKV representation, yielding the same complex modulus $\mathbf{G}^{*}(\omega)$, in accordance with the equivalence stated in Eq. (10). Equation (69) highlights that GM and GKV elements naturally generate a discrete relaxation spectrum, with characteristic relaxation time matrices $\bm{\tau}_{i}=\bar{\mathbf{C}}_{i}\bar{\mathbf{K}}_{i}^{-1}=\mathbf{C}_{i}\mathbf{K}_{i}^{-1}$, $i\in\{1,\dots,n\}$.
@@ A.3 FrBDn+1 model parameters and interconversions
Once the stiffness and damping matrices $\mathbf{K}_{i}$ and $\mathbf{C}_{i}$ have been identified, they can be directly employed in the FrBDn+1 models introduced in Sect. 2.2 by introducing the normalised parameters $\bar{\mathbf{K}}_{i}=\mathbf{K}_{i}/p$, $i\in\{0,\dots,n\}$, and $\bar{\mathbf{C}}_{i}=\mathbf{C}_{i}/p$, $i\in\{1,\dots,n\}$. where $p(\bm{x},s)$ denotes the local normal pressure. For the GM realisation, these parameters enter the dynamics through Eq. (19), whereas for the GKV realisation they appear in Eq. (20).